OpenStudy (anonymous):

a. Write functions for the following transformations using function notation. Choose a different letter to represent each function. For example, you can use R to represent rotations. Assume that a positive rotation occurs in the counterclockwise direction. • translation of a units to the right and b units up • reflection across the y-axis • reflection across the x-axis • rotation of 90 degrees counterclockwise about the origin, point O • rotation of 180 degrees counterclockwise about the origin, point O • rotation of 270 degrees counterclockwise about the origin, point O

OpenStudy (anonymous):

@amorfide

OpenStudy (amorfide):

• translation of a units to the right and b units up this means you are going A in the positive x direction so if I had f(x)=x translating a in the positive X direction means I get F(x+a)= (x-a) translating b units up means you are going b in the positive Y direction so if I had f(x)=x f(x)+b = x+b so to get translation a in positive x and b in positive y I get f(x)=x f(x+a)+b=(x-a)+b • reflection across the y-axis to reflect a function across the y axis f(x)=x f(-x)=(-x) • reflection across the x-axis reflection across x axis f(x)=x -f(x)= -(x)

OpenStudy (amorfide):

I don't know how to describe rotations

OpenStudy (amorfide):

@pooja195

pooja195 (pooja195):

@TheSmartOne

OpenStudy (anonymous):

@amorfide thank you so much, you're a life saver. I just need to answer 1 more and i'm finished with my entire online class lol

OpenStudy (anonymous):

@amorfide c. In the same way that other functions can be combined, a series of transformations can be combined into a single function. For example, this statement for the function S shows one way to represent the rotation of a point 270° counterclockwise about the origin followed by a translation 3 units to the left and 1 unit up: S(270, O)/<-3,1>(x, y) = (y – 3, -x + 1). Write a function to represent each series of transformations: • rotation of 90 degrees counterclockwise about the origin, point O, then a reflection across the x-axis • reflection across the y-axis, then a translation a units to the right and b units up • translation a units to the right and b units up, then a rotation of 180 degrees counterclockwise about the origin, then a reflection across the y-axis

OpenStudy (amorfide):

I can not do rotations I am sorry post this as a question and hope someone comes along

OpenStudy (anonymous):

@amorfide its fine but thank you so much !

OpenStudy (amorfide):

np I try to help :)